a particle of mass m is moving in a horizontal circle of radius r with constant angular speed w the areal velocity of the particle is 

Grade:11

1 Answers

Vikas TU
4 years ago
Dear student  
Centripetal force = mv^2/r = k/r^2
Kinetic energy = 1/2 mv^2
Potential energy = - Integral F.dr 
Put the limit . from minus infinity to r .
Apply integral 
Potential Energy = -k/r 
Total Energy = k/2r - k/r = -k/2r